# xCOMET-lite: Bridging the Gap Between Efficiency and Quality in Learned MT Evaluation Metrics

### v1

@@ 5 Discussion
disable]SE: what is the lesson learned? Which of the three approaches is best? Can they be combined? This should be in a real discussion section.
@@
Pruning methods, while capable of reducing memory consumption and increasing throughput, result in a more noticeable decrease in correlation compared to quantization. Our results align with the findings in Rei et al. (2022b), which conclude that up to 5 out of 24 layers of encoder model can be removed without noticeable quality degradation of the metric. At the same time, the layer pruning works slightly worse than in other tasks (Gromov et al., 2024; Men et al., 2024), disable]SE: same or different tasks?; DL: different where up to 50% of layers could be removed for large models. Pruning appears incompatible with our distilled model, due to a substantial drop in metric quality.
@@ 6 Conclusion
In the rapidly evolving field of MT evaluation, the current top-performing metrics, such as MetricX, xCOMET, and GEMBA-MQM, are all based on extremely large underlying models. These models, including mT5 with 13B parameters, XLM-RoBERTa-XXL with 11B parameters, and the closed-source GPT-4 with an estimated 1.7T parameters, pushed the boundaries of performance but come with substantial computational costs and hardware requirements. Our research aims to address these challenges by comparing three commonly used compression methods  — quantization, pruning, and knowledge distillation  — in compressing the xCOMET model. We have demonstrated that these methods can effectively reduce memory consumption and increase processing speed while maintaining competitive performance, making them viable options for deploying large state-of-the-art learned metric for MT evaluation in a resource-constrained environments. In particular, our distilled model xCOMET-lite achieves competitive prediction quality with a substantially smaller model size, offering a solution for researchers and practitioners with no access to top-tier hardware.
Based on our findings, we recommend the following: for the highest quality with a reduced VRAM requirements, opt for 8-bit or 3-bit quantization with GPTQ. For improved speed without substantial quality penalty, test 4-bit quantization with QLoRA or prune up to 25% of the model layers. For massive speedup and low hardware requirements, consider the distilled model xCOMET-lite or its quantized version, accepting a slight compromise on quality. The choice of compression method ultimately depends on the hardware, amount of data, and acceptable quality loss. disable]SE: the ending is good; DL: yay!
@@ 7 Limitations
  • Our study focuses solely on machine translation evaluation and does not consider other tasks, disable]SE: do we have reason to suspect that results would not be much different? That would be important to mitigate the limitation such as summarization evaluation. To the best of our knowledge, all currently existing summarization evaluation metrics are regression-only and do not offer error span prediction. Therefore, it is unclear if the results would be different for this task. Future research could explore the applicability of these compression methods to a broader range of natural language processing tasks.
  • Our measure of a metric quality, Kendall-$\tau$ correlation with human judgments, is known to incorrectly reward metrics for predicting ties (Deutsch et al., 2023). disable]SE: all models suffer equally?; DL: every model that uses kendall as it’s measure of quality i suppose
@@ Appendix F Results on Eval4NLP dataset
Table 6: Summary table for different compression methods in setting without reference translations. Average is computed over three language pairs for Kendall correlation. For peak memory the mean and maximum values are computed, and the maximum is reported in parentheses. XL stands for xCOMET-XL, XXL – xCOMET-XXL.

### v2

@@ 5 Discussion
Pruning methods, while capable of reducing memory consumption and increasing throughput, result in a more noticeable decrease in correlation compared to quantization. Our results align with the findings in Rei et al. (2022b), which conclude that up to 5 out of 24 layers of encoder model can be removed without noticeable quality degradation of the metric. At the same time, the layer pruning works slightly worse than in other tasks (Gromov et al., 2024; Men et al., 2024), where up to 50% of layers could be removed for large models. Pruning appears incompatible with our distilled model, due to a substantial drop in metric quality. Magnitude pruning with 4:8 sparsity pattern shows promising results with respect to quality / speedup trade-off. Moreover, it potentially offers almost 50% reduction in peak memory consumption (and e.g. torch library will likely support structured spasity formats quite soon).
@@ 6 Conclusion
In the rapidly evolving field of MT evaluation, the current top-performing metrics, such as MetricX, xCOMET, and GEMBA-MQM, are all based on extremely large underlying models. These models, including mT5 with 13B parameters, XLM-RoBERTa-XXL with 11B parameters, and the closed-source GPT-4 with an estimated 1.7T parameters, pushed the boundaries of performance but come with substantial computational costs and hardware requirements.
Our research aims to address these challenges by comparing three commonly used compression methods  — quantization, pruning, and knowledge distillation  — in compressing the xCOMET model. We have demonstrated that these methods can effectively reduce memory consumption and increase processing speed while maintaining competitive performance, making them viable options for deploying large state-of-the-art learned metric for MT evaluation in a resource-constrained environments. In particular, our distilled model xCOMET-lite achieves competitive prediction quality with a substantially smaller model size, offering a solution for researchers and practitioners with no access to top-tier hardware.
Based on our findings, we recommend the following: for the highest quality with a reduced VRAM requirements, opt for 8-bit or 3-bit quantization with GPTQ. For improved speed without substantial quality penalty, test 4-bit quantization with QLoRA, try structured magnitude pruning (2:4, 4:8) or prune up to 25% of the model layers.For massive speedup and low hardware requirements, consider the distilled model xCOMET-lite or its quantized version, accepting a slight compromise on quality. The choice of compression method ultimately depends on the hardware, amount of data, and acceptable quality loss.
Acknowledgments
The NLLG group gratefully acknowledges support from the Federal Ministry of Education and Research (BMBF) via the research grant “Metrics4NLG” and the German Research Foundation (DFG) via the Heisenberg Grant EG 375/5-1.
The NLLG group acknowledges support by the state of Baden-Württemberg through bwHPC.
@@ 7 Limitations
  • Our study focuses solely on machine translation evaluation and does not consider other tasks, such as summarization evaluation. To the best of our knowledge, all currently existing summarization evaluation metrics are regression-only and do not offer error span prediction. Therefore, it is unclear if the results would be different for this task. Future research could explore the applicability of these compression methods to a broader range of natural language processing tasks.
  • Our measure of a metric quality, Kendall-$\tau$ correlation with human judgments, is known to incorrectly reward metrics for predicting ties (Deutsch et al., 2023).
@@ Appendix F Results on Eval4NLP dataset
Table 6: An overview table with some representative results for various compression methods in setting without reference translations. Average is computed over three language pairs for Kendall correlation. For peak memory the mean and maximum values are computed, and the maximum is reported in parentheses. XL stands for xCOMET-XL, XXL – xCOMET-XXL.